. How much energy is stored in a 150µF capacitor if a 12V battery is connected across




it? How much charge is stored in this capacitor?

1
Expert's answer
2022-03-06T18:12:37-0500

Charge

"Q=CV"


"Q=12\times150\times10^{-6}=18\times10^{-4}C"

Energy

"U=\frac{1}{2}CV^2"

"U=\frac{1}{2}\times{150\times10^{-6}\times12^2}=10.8\times10^{-3}J"